An inlet pipe can fill a tank in 8 minutes. A drain can empty the tank in 12 mintes. If the tank is empty and both the pipe and drain are open, how long will it take before the tank overflows? Thanks!

Answer by edjones(8007)

Let t = time in minutes
1/8 - 1/12 = 1/t
3/24 - 2/24 = 1/24
1/24 of the job of filling the tank is done in one minute
24 min to fill the tank.
.
Ed